OpenStudy (valirium):

What is the expression in factored form? x^2 + 13x + 42

OpenStudy (tkhunny):

Have you considered searching for the factors? 42 = 1 * 41 42 = 2 * 21 Are there others? Can you find a pair that adds up to 13?

OpenStudy (vuriffy):

@Valirium Did you figure out the values yet?

OpenStudy (valirium):

There isn't any? @tkhunny @Vuriffy

OpenStudy (valirium):

for 13, I mean

OpenStudy (tkhunny):

What other factors did you find?

OpenStudy (tkhunny):

42 = 3 * 14 More?

OpenStudy (valirium):

6 * 7 for 42

OpenStudy (vuriffy):

Try values that add up to 13. 4 9 5 8 6 7 Then multiply them to see if they create 42.

OpenStudy (vuriffy):

Now, figure out if both shall be positive when expanded to add to 13.

OpenStudy (vuriffy):

(x+6)(x-7) Would not be correct, because you would receive x^2 - x - 42. Try (x+6)(x+7) for me?

OpenStudy (tkhunny):